Alex Mason has a wide-curving, uphill driveway leading to his garage. When there is a heavy snow, Alex hires a local carpenter, who shovels snow on the side in the winter, to shovel his driveway. The snow shoveler charges $30 to shovel the driveway. Following is a probability distribution of the number of heavy snows each winter.


Heavy Snows


Probability


1


1.12


2


0.19


3


0.24


4


0.22


5


0.13


Alex is considering the purchase of a new self-propelled snow blower for $575 that would allow him, his wife, or his children to clear the driveway after a snow. Discuss what you think Alex's decision should be and why.


1
Expert's answer
2021-06-24T03:58:19-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c}
 Heavy\ Snows & Probability \\ \hline
 1 & 0.12\\
 2 & 0.19\\
 3 & 0.24\\
 4 & 0.22\\
 5 & 0.13\\
 6 & 0.08\\
 7 & 0.02\\
\hdashline
 & 1.00\\
\end{array}"

What is the annual expected cost of shoveling snow?


"30(1(0.12)+2(0.19)+3(0.24)+4(0.22)"

"+5(0.13)+6(0.08)+7(0.02))=101.10"

How many winters would it take to completely payback the purchase of a snow blower?


"\dfrac{575}{101.10}=6"

6 years would it take to completely payback the purchase of a snow blower.